# Quantifying task-relevant representational similarity using decision variable correlation

### v3

@@ 4.5 Comparison to error consistency based on Cohen’s Kappa
One method that is highly relevant to DVC is Cohen’s Kappa. As a classic statistical measure of inter-rater consistency [cohenCoefficientAgreementNominal1960], Cohen’s Kappa was recently applied to quantify the error consistency of deep networks and brains [geirhosAccuracyQuantifyingTrialbytrial2020, geirhosPartialSuccessClosing2021]. These studies examined human-model alignment using categorical judgements on assorted out-of-distribution stimuli. For example, ‘edges’ where only the edges are illustrated, or ‘silhouettes’ which are filled outlines of objects. These studies arrived at very different conclusions, namely that (i) model-model similarity is significantly higher than human-human similarity, (ii) model-human similarity is extremely low and that (iii) models trained on rich datasets are more aligned with humans. At a high level, these results seem to be inconsistent with our findings, because we found that (i) deep networks exhibit modest consistency with the brain; (ii) DVCs between different deep networks trained on ImageNet-1k are not exceedingly large; (iii) models trained on rich datasets have lower DVCs with brains.
@@
Accuracy difference While Cohen’s Kappa was originally proposed to disentangle the accuracy and consistency of the observers under certain conditions, it remains possible that the two are intermingled in practice [geirhosAccuracyQuantifyingTrialbytrial2020]. To assess this issue, we examine the behavioral responses to ‘edges’ stimuli from  [geirhosImageNettrainedCNNsAre2018]. From the results shown in Fig. 5d,e, it is evident that Cohen’s Kappa is strongly affected by the difference in the accuracy of the two observers [^4] Theoretical upper-bound is given by $\kappa\;\leq\;\frac{(1-d)^{2}}{1+d^{2}}$, where $d$ is the accuracy difference between the two observers. See Appendix A.2 for the derivation.. When the difference in accuracy is high, Cohen’s kappa is low. This suggests that the low error consistency between humans and models reported in  [geirhosAccuracyQuantifyingTrialbytrial2020] is at least partially due to the large differences in accuracy between the two observers. It also potentially explains why models trained on ‘rich’ datasets, which enjoy better out-of-distribution performance, also exhibit better alignment with human behavior based on Cohen’s Kappa [geirhosPartialSuccessClosing2021].
@@
Decoder bias According to signal detection theory, Cohen’s Kappa is determined by both the correlation of the decision variables and decision criterion. Thus, we wondered if the extremely high Cohen’s Kappa values between different networks as reported in [geirhosAccuracyQuantifyingTrialbytrial2020] are due to biases in the decisions. We thus perform a second analysis to further investigate the data from  [geirhosImageNettrainedCNNsAre2018] and analysis used in  [geirhosAccuracyQuantifyingTrialbytrial2020]. Consistent with our hypothesis, we find that the approach in  [geirhosAccuracyQuantifyingTrialbytrial2020] introduces high decision bias (see Fig. 5f) and reduced accuracy, especially when target categories do not align cleanly with the original training labels. We also find that the origin of this large decision bias is because the analysis in [geirhosAccuracyQuantifyingTrialbytrial2020] is based on an aggregated decoder that estimates class probabilities by combining probabilities from related ImageNet-1k classes. Once we substituted the original decoder with a cross-validated logisitic regression classifier, the estimated Cohen’s Kappa values become largely consistent with the DVC we obtained on the main dataset we analyzed. These results suggest that the large Cohen’s Kappa values between different networks are due to the biases in the decoders [^5] See Appendix D.3 for results on the ‘edges’ data using a logistic regression classifier..
@@
These results highlight the advantages of the DVC method. DVC is insensitive to the decision biases, while the error consistency quantified based on Cohen’s Kappa captures both shared behavior biases and consistency in their underlying decision variables. This point is further demonstrated using a simple simulation (see Fig. 5e). Here, we add behavioral biases to the two observers by shifting the decision criteria consistently, so that they both prefer one category over another. Naturally, as the bias increases, the accuracy drops. Cohen’s Kappa strongly reflects these decision biases, whereas DVC remains consistent (Fig.5g). Furthermore, DVC can effectively decouple accuracy v.s. trial-by-trial consistency of the decision variables [sebastianDecisionvariableCorrelation2018].
@@ 5 Discussion
We have developed a method, DVC, to quantify the consistency of two neural representations. The focus on task-relevant features makes DVC different from other popular approaches such as RSA[kriegeskorteRepresentationalSimilarityAnalysis2008a] or linear regression [yaminsUsingGoaldrivenDeep2016]. Two representations may have high DVC yet low consistency according to RSA, or vice versa. For behavioral metrics that aim to characterize trial-by-trial consistency, one challenge has been to decouple task performance and trial-by-trial consistency. DVC provides a principled way to do so. For future work, it would be interesting to combine the analysis of DVC at the behavioral level [sebastianDecisionvariableCorrelation2018] and the neural representation level to dissect the contribution of consistency of DVs and the shared biases of the observers. It would also be interesting to systematically compare DVC to other proposed similarity measures [kriegeskorteRepresentationalSimilarityAnalysis2008a, yaminsUsingGoaldrivenDeep2016, schrimpfBrainScoreWhichArtificial2020, williamsGeneralizedShapeMetrics2022a] [^6] See Appendix E for some preliminary results on comparison to RSA.. Applying the DVC method, we find that there are surprising negative correlations between (i) the classification performance of the deep networks trained on ImageNet-1k and (ii) the consistency with the neural representation in V4/IT. Furthermore, training the deep network adversarially or using rich datasets seems to evoke decrease, rather than increase of DVCs. While it is unclear how to close the gap between the image-by-image consistency of deep networks to that of the brain, the following directions might be promising: (i) training networks using datasets that better resemble the stimulus statistics that drives the evolution of the primate visual system [mehrerEcologicallyMotivatedImage2021]; (ii) develop training procedures that better capture the stimulus noise and internal noise of the brain [dapelloSimulatingPrimaryVisual2020], as well as low level properties of the visual system (e.g., optics and foveation).
@@
Acknowledgment We thank Nikolaus Kriegeskorte for helpful discussions. This research was supported by the National Institute Of Neurological Disorders And Stroke of the National Institutes of Health (Award Number R01NS133924) and a Sloan Research Fellowship to X.-X. Wei. The content is solely the responsibility of the authors and does not necessarily represent the official views of the funding agencies.
@@ C.4 DVC results are consistent in another neural datasets
To alleviate concern that the dataset from which we derived the main results contains only two monkeys, we verified the result on an additional dataset from Bashivan et al [Bashivan2019neural]. The dataset contains three monkey subjects viewing natural and synthesized images. Among the three subjects, only one (’monkey M’) has sufficient neuron count for this test (n = 168). The neurons are pooled across four recording sessions. We used data from monkey M viewing naturalistic images and calculated its DVC against neural networks viewing the same images. The main results are also significant in this case (Pearson’s correlation = -0.84, p = 0.87e-05, Fig.A.4a).
Figure A.4: The main result is confirmed in an additional dataset from [Bashivan2019neural]. (a) DVC of monkey M against representative models. Again, these is a strong negative correlation between the top-1 accuracy of the deep networks and the DVC to the monkey brain. (b) DVC between representative models.
@@ Appendix D Additional results on the effect of behavioral decoding on Cohen’s Kappa
Geirhos et al. [geirhosAccuracyQuantifyingTrialbytrial2020] have discussed several caveats of applying Cohen’s Kappa, including that Cohen’s Kappa is bounded by error overlap expected by chance $c_{exp}$. We also explicitly showed that the accuracy difference $d=|p_{i}-p_{j}|$ provides an upper bound on Cohen’s Kappa (see Section A.2). When the accuracy difference is high, one subject is often right, while the other is often wrong, then their behaviors cannot be consistent. From the original behavioral data, it is clear that human subjects perform with high accuracy but models perform poorly(Fig. A.6). While Geirhos et al used simulations to show that under the condition that the subjects act independently, accuracy is not necessarily correlated with model performance, Cohen’s Kappa may still depend on accuracy under more general conditions.
Figure A.6: Re-plotting of the data (based on "edge" stimuli from [geirhosAccuracyQuantifyingTrialbytrial2020]. Red: accuracy of human subjects. Blue: accuracy of deep network models tested. Overall, the accuracy of the deep network models are much lower than that of human subjects.
@@ Appendix E Comparison to Representational Similarity Analysis
Representational similarity analysis (RSA) measures the similarity between the general arrangement of categories (or samples) between two representations. Under appropriate conditions, it can be seen as equivalent to centered kernel alignment (CKA) and canonical correlation analysis (CCA) [williamsEquivalenceRepresentationalSimilarity2024]. In the monkey dataset that we tested, RSA is positively correlated with DVC when measuring both model-monkey alignments and model-model alignments (see Fig. A.8a-d)).

### v4

@@ 4.5 Comparison to error consistency based on Cohen’s Kappa
One method that is highly relevant to DVC is Cohen’s Kappa. As a classic statistical measure of inter-rater consistency [6], Cohen’s Kappa was recently applied to quantify the error consistency of deep networks and brains [16, 17]. These studies examined human-model alignment using categorical judgements on assorted out-of-distribution stimuli. For example, ‘edges’ where only the edges are illustrated, or ‘silhouettes’ which are filled outlines of objects. These studies arrived at very different conclusions, namely that (i) model-model similarity is significantly higher than human-human similarity, (ii) model-human similarity is extremely low and that (iii) models trained on rich datasets are more aligned with humans. At a high level, these results seem to be inconsistent with our findings, because we found that (i) deep networks exhibit modest consistency with the brain; (ii) DVCs between different deep networks trained on ImageNet-1k are not exceedingly large; (iii) models trained on rich datasets have lower DVCs with brains.
@@
Accuracy difference While Cohen’s Kappa was originally proposed to disentangle the accuracy and consistency of the observers under certain conditions, it remains possible that the two are intermingled in practice [16]. To assess this issue, we examine the behavioral responses to ‘edges’ stimuli from  [18]. From the results shown in Fig. 5d,e, it is evident that Cohen’s Kappa is strongly affected by the difference in the accuracy of the two observers [^4] Theoretical upper-bound is given by $\kappa\;\leq\;\frac{(1-d)^{2}}{1+d^{2}}$, where $d$ is the accuracy difference between the two observers. See Appendix A.2 for the derivation.. When the difference in accuracy is high, Cohen’s kappa is low. This suggests that the low error consistency between humans and models reported in  [16] is at least partially due to the large differences in accuracy between the two observers. It also potentially explains why models trained on ‘rich’ datasets, which enjoy better out-of-distribution performance, also exhibit better alignment with human behavior based on Cohen’s Kappa [17].
@@
Decoder bias According to signal detection theory, Cohen’s Kappa is determined by both the correlation of the decision variables and decision criterion. Thus, we wondered if the extremely high Cohen’s Kappa values between different networks as reported in [16] are due to biases in the decisions. We thus perform a second analysis to further investigate the data from  [18] and analysis used in  [16]. Consistent with our hypothesis, we find that the approach in  [16] introduces high decision bias (see Fig. 5f) and reduced accuracy, especially when target categories do not align cleanly with the original training labels. We also find that the origin of this large decision bias is because the analysis in [16] is based on an aggregated decoder that estimates class probabilities by combining probabilities from related ImageNet-1k classes. Once we substituted the original decoder with a cross-validated logisitic regression classifier, the estimated Cohen’s Kappa values become largely consistent with the DVC we obtained on the main dataset we analyzed. These results suggest that the large Cohen’s Kappa values between different networks are due to the biases in the decoders [^5] See Appendix D.3 for results on the ‘edges’ data using a logistic regression classifier..
@@
These results highlight the advantages of the DVC method. DVC is insensitive to the decision biases, while the error consistency quantified based on Cohen’s Kappa captures both shared behavior biases and consistency in their underlying decision variables. This point is further demonstrated using a simple simulation (see Fig. 5e). Here, we add behavioral biases to the two observers by shifting the decision criteria consistently, so that they both prefer one category over another. Naturally, as the bias increases, the accuracy drops. Cohen’s Kappa strongly reflects these decision biases, whereas DVC remains consistent (Fig.5g). Furthermore, DVC can effectively decouple accuracy v.s. trial-by-trial consistency of the decision variables [39].
@@ 5 Discussion
We have developed a method, DVC, to quantify the consistency of two neural representations. The focus on task-relevant features makes DVC different from other popular approaches such as RSA[25] or linear regression [52]. Two representations may have high DVC yet low consistency according to RSA, or vice versa. For behavioral metrics that aim to characterize trial-by-trial consistency, one challenge has been to decouple task performance and trial-by-trial consistency. DVC provides a principled way to do so. For future work, it would be interesting to combine the analysis of DVC at the behavioral level [39] and the neural representation level to dissect the contribution of consistency of DVs and the shared biases of the observers. It would also be interesting to systematically compare DVC to other proposed similarity measures [25, 52, 38, 49] [^6] See Appendix E for some preliminary results on comparison to RSA.. Applying the DVC method, we find that there are surprising negative correlations between (i) the classification performance of the deep networks trained on ImageNet-1k and (ii) the consistency with the neural representation in V4/IT. Furthermore, training the deep network adversarially or using rich datasets seems to evoke decrease, rather than increase of DVCs. While it is unclear how to close the gap between the image-by-image consistency of deep networks to that of the brain, the following directions might be promising: (i) training networks using datasets that better resemble the stimulus statistics that drives the evolution of the primate visual system [31]; (ii) develop training procedures that better capture the stimulus noise and internal noise of the brain [10], as well as low level properties of the visual system (e.g., optics and foveation).
@@
Acknowledgments We thank Nikolaus Kriegeskorte for helpful discussions. This research was supported by the National Institute Of Neurological Disorders And Stroke of the National Institutes of Health (Award Number R01NS133924), the NIH BRAIN Initiative and National Institute On Drug Abuse (Award Number R01DA060742), and a Sloan Research Fellowship (to X.-X. Wei). The content is solely the responsibility of the authors and does not necessarily represent the official views of the funding agencies.
@@ Appendix C Implementation and verification
For detailed implementation used in analysis, refer to https://github.com/wei-bbc-lab/DVC.
@@ C.4 DVC results are consistent in another neural datasets
To alleviate concern that the dataset from which we derived the main results contains only two monkeys, we verified the result on an additional dataset from Bashivan et al [2]. The dataset contains three monkey subjects viewing natural and synthesized images. Among the three subjects, only one (’monkey M’) has sufficient neuron count for this test (n = 168). The neurons are pooled across four recording sessions. We used data from monkey M viewing naturalistic images and calculated its DVC against neural networks viewing the same images. The main results are also significant in this case (Pearson’s correlation = -0.84, p = 0.87e-05, Fig.A.4a).
Figure A.4: The main result is confirmed in an additional dataset from [2]. (a) DVC of monkey M against representative models. Again, these is a strong negative correlation between the top-1 accuracy of the deep networks and the DVC to the monkey brain. (b) DVC between representative models.
@@ Appendix D Additional results on the effect of behavioral decoding on Cohen’s Kappa
Geirhos et al. [16] have discussed several caveats of applying Cohen’s Kappa, including that Cohen’s Kappa is bounded by error overlap expected by chance $c_{exp}$. We also explicitly showed that the accuracy difference $d=|p_{i}-p_{j}|$ provides an upper bound on Cohen’s Kappa (see Section A.2). When the accuracy difference is high, one subject is often right, while the other is often wrong, then their behaviors cannot be consistent. From the original behavioral data, it is clear that human subjects perform with high accuracy but models perform poorly(Fig. A.6). While Geirhos et al used simulations to show that under the condition that the subjects act independently, accuracy is not necessarily correlated with model performance, Cohen’s Kappa may still depend on accuracy under more general conditions.
Figure A.6: Re-plotting of the data (based on "edge" stimuli from [16]. Red: accuracy of human subjects. Blue: accuracy of deep network models tested. Overall, the accuracy of the deep network models are much lower than that of human subjects.
@@ Appendix E Comparison to Representational Similarity Analysis
Representational similarity analysis (RSA) measures the similarity between the general arrangement of categories (or samples) between two representations. Under appropriate conditions, it can be seen as equivalent to centered kernel alignment (CKA) and canonical correlation analysis (CCA) [50]. In the monkey dataset that we tested, RSA is positively correlated with DVC when measuring both model-monkey alignments and model-model alignments (see Fig. A.8a-d)).